### v1

@@ 2.4 Holomorphic regularity
Consider $f:\mathcal{O}\to\mathbb{C}$ where $\mathcal{O}\subseteq\mathbb{C}^{d}$ and $d\in\mathbb{N}$. The function $f$ is holomorphic in $\mathcal{O}$ if the following limit exists for every $\boldsymbol{z}\in\mathcal{O}$ and every $j\in[d]$:
$\lim_{\begin{subarray}{c}h\in\mathbb{C}\\ h\to 0\end{subarray}}\frac{f(\boldsymbol{z}+h\boldsymbol{e}_{j})-f(\boldsymbol{z})}{h},$
where $\boldsymbol{e}_{j}$ is the $j$-th element of the canonical basis of $\mathbb{R}^{d}$.
@@
The filled-in Berstein polyellipse of parameter $\boldsymbol{\rho}=(\rho_{j})_{j=1}^{d}\in\mathbb{R}^{d}$, with $\boldsymbol{\rho}>\boldsymbol{1}$ (that is, $\rho_{j}>1$ for $j=1,2,...,d$), denoted as $\mathcal{E}_{\boldsymbol{\rho}}$, is the Cartesian product of filled-in Berstein ellipses of parameters $\rho_{j}$, defined as
$\mathcal{E}_{\rho_{j}}=\left\{\frac{z+z^{-1}}{2}:z\in\mathbb{C},1\leq\lvert z\rvert\leq\rho_{j}\right\}.$
Namely,
$\mathcal{E}_{\boldsymbol{\rho}}=\mathcal{E}_{\rho_{1}}\times\mathcal{E}_{\rho_{2}}\times...\times\mathcal{E}_{\rho_{d}}.$
Assume that $f\in L_{\varrho}^{2}(\mathcal{U})$ can be extended in a holomorphic way to an open set $\mathcal{O}\subseteq\mathbb{C}^{d}$ such that $\mathcal{E}_{\rho}\subset\mathcal{O}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$, i.e., there exists $\tilde{f}:\mathcal{O}\to\mathbb{C}$ holomorphic such that $\tilde{f}|_{\mathcal{U}}=f$. Then, the best $s$-term approximation $\sigma_{s}(\boldsymbol{c})_{q}$ decays algebraically fast in $s$, with dimension-independent decay rate [5, Theorem 3.6]. Namely,
(15) $\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{\lVert f\rVert_{L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})}\cdot C(d,p,\boldsymbol{\rho})}{(s+1)^{\frac{1}{p}-\frac{1}{q}}},$
for all $0<p\leq q\leq\infty$ and $s\in\mathbb{N}_{0}$, where $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is the infinite vector of coefficients of $f$, $C(d,p,\boldsymbol{\rho})>0$ is a constant depending on $d$, $p$ and $\boldsymbol{\rho}$ explicitly defined in [5, Theorem 3.6], and $L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})$ is the space of essentially bounded functions $f:\mathcal{E}_{\boldsymbol{\rho}}\to\mathbb{C}$.
Furthermore, with the holomorphic assumption on $f$, it si possible to show quasi-optimal convergence guarantees of sparse polynomial approximation via least squares or compressed sensing. Indeed, denoting with $\hat{f}_{\mathsf{LS}}$ a solution to problem (9) for a suitable polynomial space $S$, we have
@@
with probability at least $1-\epsilon$; see [2, Theorem 6.1] and Appendix B. Similarly, denoting with $\hat{f}_{\mathsf{CS}}$ a solution to problem (10) respectively, we have
@@
with probability at least $1-\epsilon$ and where
$\tilde{m}=\frac{m}{c\cdot\log(2m)\cdot[\log(2m)\cdot\min\{\log(m)+d,\log(2d)\cdot\log(2m)\}+\log(\epsilon^{-1})]},$
for some universal constant $c>0$. Consequently, least squares and compressed sensing approximation achieve the same algebraic rate of convergence as that of the best $s$-term approximation in terms of $m$, up to log factors [5, Theorem 7.12].
@@ 2.5 Community structure assumption for diffusion on graphs
Consider a weighted graph $G=(V,E)$ with weighted adjacency matrix $\boldsymbol{W}$ of bounded size (i.e., $\lvert V\rvert<\infty$) and with $K$ communities $\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{K}$. In the time-and-edge-dependent diffusion case, we consider the parametric map
(16) $\boldsymbol{C}:[0,T]\times\mathcal{U}\to\mathbb{C}^{|V|\times|V|}$
@@
In our setup, we assign the same diffusion coefficient to the set of edges linking a pair of communities ($\mathcal{C}_{i}$, $\mathcal{C}_{j}$), for $i,j=1,2,...,K$. Thus, for a parameter $\boldsymbol{y}=(y_{1},...,y_{d})$, we assign one entry $y_{k}$ of the parameter to the set of edges between the communities denoted by the pair ($\mathcal{C}_{i}$, $\mathcal{C}_{j}$) and, by symmetry, the pair ($\mathcal{C}_{j}$, $\mathcal{C}_{i}$). As a result, we have the relation $d=\frac{K(K+1)}{2}$ between the dimension $d$ and the number of communities $K$. The resulting image is a symmetric block matrix $\boldsymbol{C}(t)$, where each submatrix $C_{ij}$ (or $C(t)_{ij}$) gives the diffusion coefficients along all the edges between the communities $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$.
@@
The following Assumption, that will be recalled in the main result of this paper (Theorem 3.1), summarizes the requirements just discussed.
@@
There exist a compact set $\mathcal{K}\subset\mathbb{C}^{d}$, with $\mathcal{U}\subset\mathcal{K}$, and a map $\tilde{\boldsymbol{C}}:[0,T]\times\mathcal{K}\to\mathbb{C}^{|V|\times|V|}$ such that $\tilde{\boldsymbol{C}}|_{[0,T]\times\mathcal{U}}=\boldsymbol{C}$, $\tilde{\boldsymbol{C}}$ is continuous in $[0,T]\times\mathcal{K}$ and $\tilde{\boldsymbol{C}}(t,\cdot)$ is holomorphic in $\mathring{\mathcal{K}}$ for every $t\in[0,T]$. Moreover, we consider $\boldsymbol{C}$ of the form
@@
where $\boldsymbol{C}^{(i,j)}(t,\boldsymbol{y})\in\mathbb{R}^{|\mathcal{C}_{i}|\times|\mathcal{C}_{j}|}$ satisfies $\boldsymbol{C}^{(i,j)}(t,\boldsymbol{y})=c^{(i,j)}(t,\boldsymbol{y})\cdot\mathds{1}$ with $\mathds{1}$ being an $|\mathcal{C}_{i}|\times|\mathcal{C}_{j}|$ matrix of 1’s and such that $c^{(i,j)}(t,\boldsymbol{y})=c^{(j,i)}(t,\boldsymbol{y})$ for all $i,j\in[K]$. In particular, $\boldsymbol{C}(t,\boldsymbol{y})$ is symmetric. The number of distinct functions $c^{(i,j)}(t,\boldsymbol{y})$ is $d=\frac{K(K+1)}{2}$.
@@ 3 Theoretical guarantees
(17) $f(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\cdot,\boldsymbol{y}))$
for a given node $v$, where $\boldsymbol{u}(\cdot,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ is the unique solution to the problem (4) with $\boldsymbol{C}(t)=\boldsymbol{C}(t,\boldsymbol{y})$. As illustrated in Section 2.4, functions that can be extended in a holomorphic way have desirable properties with respect to the recoverability of their polynomial approximation; hence our goal is to show that there exists a holomorphic extension for $f$. We will use [40] and [5, Lemma 4.3] which states the following:
@@
(18) $\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert\leq L\lVert\boldsymbol{u}-\boldsymbol{v}\rVert,\quad\forall t,s\in[0,T],\quad\forall\boldsymbol{u},\boldsymbol{v}\in\mathbb{C}^{n},\forall\boldsymbol{z}\in\mathcal{K},$
[…]
admits a unique solution $t\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ for every $\boldsymbol{z}\in\mathcal{K}$ and the mapping $(t,\boldsymbol{z})\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ is continuous in $[0,T]\times\mathcal{K}$. In addition, let $\mathring{\mathcal{K}}$ be the interior of $\mathcal{K}$ and assume that $\boldsymbol{z}\mapsto\boldsymbol{g}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for any fixed $t\in[0,T]$ and $(\boldsymbol{v},\boldsymbol{z})\mapsto\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})$ is holomorphic in $\mathbb{C}^{n}\times\mathring{\mathcal{K}}$ for any fixed $(t,s)\in[0,T]^{2}$. Then $\boldsymbol{z}\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for any fixed $t\in[0,T]$.
@@
Theorem 3.1 (Holomorphy of parametric graph diffusion).
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2 for some compact $\mathcal{K}\subseteq\mathbb{C}^{d}$ such that $\mathcal{K}\subset\mathcal{U}$. Then the map $\boldsymbol{f}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}^{|V|}$ defined by $\boldsymbol{f}(\boldsymbol{y})=\boldsymbol{u(}T,\boldsymbol{y})$ satisfying problem (4) admits a holomorphic extension $\tilde{\boldsymbol{f}}$ to $\mathring{\mathcal{K}}$. Moreover, for any compact subset $\mathcal{H}\subset\mathring{\mathcal{K}}$, we have
(19) $\|\tilde{f}\|_{L^{\infty}(\mathcal{H)}}\leq B(\boldsymbol{u}_{0},T,\boldsymbol{M}),$
@@
(20) $B(\boldsymbol{u}_{0},T,\boldsymbol{M})=\|\boldsymbol{u}_{0}\|_{2}\cdot\exp\left(2\cdot\int_{0}^{T}\sup_{\boldsymbol{z}\in\mathcal{H}}\|\boldsymbol{M}(t,\boldsymbol{z})\|_{2\to 2}\,dt\right)<\infty$
@@
We define a holomorphic extension $\tilde{\boldsymbol{f}}$ in a natural way using the complex-valued extension $\tilde{\boldsymbol{C}}$ to $\boldsymbol{C}$, whose existence is guaranteed by Assumption 2. Now, observe that we can reformulate problem (4) (where $\boldsymbol{C}$ is replaced by $\tilde{\boldsymbol{C}}$) as the following parametric Volterra integral equation:
[…]
where $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$ and $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})=\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(t,\boldsymbol{z}),\boldsymbol{W}(t))$, with $\boldsymbol{D}$ defined as in equation (3).
@@
  • $\boldsymbol{g}(t,\boldsymbol{z})\equiv\boldsymbol{u}_{0}$ is continuous in $[0,T]\times\mathcal{K}$, and $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$ is continuous in $[0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$;
@@
Firstly, $\boldsymbol{g}(t,\boldsymbol{z})\equiv\boldsymbol{u}_{0}$ is a constant function, therefore $\boldsymbol{g}$ is continuous in $[0,T]\times\mathcal{K}$ and holomorphic in $\mathring{\mathcal{K}}$ for all fixed $t\in[0,T]$.

### v2

@@ 2.4 Holomorphic regularity
Consider $f:\mathcal{O}\to\mathbb{C}$ where $\mathcal{O}\subseteq\mathbb{C}^{d}$ and $d\in\mathbb{N}$. The function $f$ is holomorphic in $\mathcal{O}$ if the following limit exists for every $\boldsymbol{z}\in\mathcal{O}$ and every $j\in[d]$: $\lim_{h\in\mathbb{C},h\to 0}\frac{f(\boldsymbol{z}+h\boldsymbol{e}_{j})-f(\boldsymbol{z})}{h},$ where $\boldsymbol{e}_{j}$ is the $j$-th element of the canonical basis of $\mathbb{R}^{d}$.
@@
The filled-in Berstein polyellipse of parameter $\boldsymbol{\rho}=(\rho_{j})_{j=1}^{d}\in\mathbb{R}^{d}$, with $\boldsymbol{\rho}>\boldsymbol{1}$ (that is, $\rho_{j}>1$ for j=1,2,,dj=1,2,...,d), denoted as $\mathcal{E}_{\boldsymbol{\rho}}$, is the Cartesian product of filled-in Berstein ellipses of parameters $\rho_{j}$, defined as $\mathcal{E}_{\rho_{j}}=\left\{\frac{z+z^{-1}}{2}:z\in\mathbb{C},1\leq\lvert z\rvert\leq\rho_{j}\right\}$. Namely, $\mathcal{E}_{\boldsymbol{\rho}}=\mathcal{E}_{\rho_{1}}\times\mathcal{E}_{\rho_{2}}\times...\times\mathcal{E}_{\rho_{d}}$.
Assume that $f\in L_{\varrho}^{2}(\mathcal{U})$ can be extended in a holomorphic way to an open set $\mathcal{O}\subseteq\mathbb{C}^{d}$ such that $\mathcal{E}_{\rho}\subset\mathcal{O}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$, i.e., there exists $\tilde{f}:\mathcal{O}\to\mathbb{C}$ holomorphic such that $\tilde{f}|_{\mathcal{U}}=f$. Then, the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ decays algebraically fast in $s$, with dimension-independent decay rate [1, Theorem 3.6]. Namely,
$\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{\lVert f\rVert_{L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})}\cdot C(d,p,\boldsymbol{\rho})}{(s+1)^{\frac{1}{p}-\frac{1}{q}}},$ (15)
for all $0<p\leq q\leq\infty$ and $s\in\mathbb{N}_{0}$, where $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is the infinite vector of coefficients of $f$, $C(d,p,\boldsymbol{\rho})>0$ is a constant depending on $d$, $p$ and $\boldsymbol{\rho}$ explicitly defined in [1, Theorem 3.6], and $L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})$ is the space of essentially bounded functions $f:\mathcal{E}_{\boldsymbol{\rho}}\to\mathbb{C}$.
Furthermore, with the holomorphic assumption on $f$, it is possible to show quasi-optimal convergence guarantees of sparse polynomial approximation via least squares or compressed sensing. Indeed, denoting with $\hat{f}_{\mathsf{LS}}$ a solution to problem (9) for a suitable polynomial space $S$, we have
@@
with probability at least $1-\epsilon$; see [26, Theorem 6.1] and Appendix B. Similarly, denoting with $\hat{f}_{\mathsf{CS}}$ a solution to problem (10) respectively, we have
@@
[…]
@@ 2.5 Community structure assumption for diffusion on graphs
Consider a weighted graph $G=(V,E)$ with weighted adjacency matrix $\boldsymbol{W}$ of bounded size (i.e., $\lvert V\rvert$$\infty$) and with $K$ communities $\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{K}$. In the time-and-edge-dependent diffusion case, we consider the parametric map
$\boldsymbol{C}:[0,T]\times\mathcal{U}\to\mathbb{C}^{|V|\times|V|}$ (16)
@@
In our setup, we assign the same diffusion coefficient to the set of edges linking a pair of communities ($\mathcal{C}_{i}$, $\mathcal{C}_{j}$), for i,j=1,2,,Ki,j=1,2,...,K. Thus, for a parameter $\boldsymbol{y}=(y_{1},...,y_{d})$, we assign one entry $y_{k}$ of the parameter to the set of edges between the communities denoted by the pair ($\mathcal{C}_{i}$, $\mathcal{C}_{j}$) and, by symmetry, the pair ($\mathcal{C}_{j}$, $\mathcal{C}_{i}$). As a result, we have the relation
$d=\frac{K(K+1)}{2}.$ (17)
linking the dimension $d$ with the number of communities $K$. The resulting image is a symmetric block matrix $\boldsymbol{C}(t)$, where each submatrix $C_{ij}$ (or $C(t)_{ij}$) gives the diffusion coefficients along all the edges between the communities $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$. To provide some intuition about the community-based structure of the diffusivity map, we can think about, e.g., consensus dynamics in a political network (e.g., a parliament). In that context it is natural to expect that information will flow differently (i.e., faster or slower) within each party or across parties.
@@
The following Assumption, that will be used in the main result of this paper (Theorem 1), summarizes the requirements just discussed.
@@
There exists a compact set $\mathcal{K}\subset\mathbb{C}^{d}$, with $\mathcal{U}\subset\mathcal{K}$, and a map $\tilde{\boldsymbol{C}}:[0,T]\times\mathcal{K}\to\mathbb{C}^{|V|\times|V|}$ such that $\tilde{\boldsymbol{C}}|_{[0,T]\times\mathcal{U}}=\boldsymbol{C}$, $\tilde{\boldsymbol{C}}$ is continuous in $0,T]\times\mathcal{K}$ and $\tilde{\boldsymbol{C}}(t,\cdot)$ is holomorphic in $\mathring{\mathcal{K}}$ for every $t\in[0,T]$. Moreover, we consider $\boldsymbol{C}$ of the form
@@
where $\boldsymbol{C}^{(i,j)}(t,\boldsymbol{y})\in\mathbb{C}^{|\mathcal{C}_{i}|\times|\mathcal{C}_{j}|}$ satisfies $\boldsymbol{C}^{(i,j)}(t,\boldsymbol{y})=c^{(i,j)}(t,\boldsymbol{y})\cdot\mathds{1}$ with $\mathds{1}$ being an $|\mathcal{C}_{i}|\times|\mathcal{C}_{j}|$ matrix of 1’s and such that $c^{(i,j)}(t,\boldsymbol{y})=c^{(j,i)}(t,\boldsymbol{y})$ for all $i,j\in[K]$. In particular, $\boldsymbol{C}(t,\boldsymbol{y})$ is symmetric. The number of distinct functions $c^{(i,j)}(t,\boldsymbol{y})$ is $d$ as in (17).
We now provide a typical example of parametrization satisfying this assumption.
Example 1.
Given a parameter $\boldsymbol{y}=(y_{1},y_{2},...,y_{d})\in[-1,1]^{d}$, with $d$ as in (17), the matrix $\boldsymbol{C}(t,\boldsymbol{y})$ in equation (4) can be constructed using an affine parameterization, e.g.,
$\boldsymbol{C}(t,\boldsymbol{y})=\begin{pmatrix}\frac{y_{\sigma(1,1)}+1}{2}\cdot h_{\sigma(1,1)}(t)\mathds{1}&\cdots&\frac{y_{\sigma(1,K)}+1}{2}\cdot h_{\sigma(1,K)}(t)\mathds{1}\\ \vdots&\ddots&\vdots\\ \frac{y_{\sigma(K,1)}+1}{2}\cdot h_{\sigma(K,1)}(t)\mathds{1}&\cdots&\frac{y_{\sigma(K,K)}+1}{2}\cdot h_{\sigma(K,K)}(t)\mathds{1}\\ \end{pmatrix}$, (18)
where in the $(i,j)$-th block, $\mathds{1}\in\mathbb{R}^{\lvert\mathcal{C}_{i}\rvert\times\lvert\mathcal{C}_{j}\rvert}$ denotes a matrix of ones of suitable dimension, $\sigma:[K]\times[K]\mapsto[d]$ is a function that defines a unique ordering (e.g., lexicographic) on $\{(i,j)\in[K]^{2}:i\leq j\}$ and with $\sigma(i,j)=\sigma(j,i)$ for $i>j$. The functions $h_{1}(t),...,h_{d}(t)$ are continuous on $[0,T]$. Affine parameterizations are very popular in modeling parametric PDEs [1, 2] and represent a simple way to model diffusivity parameters ranging in some prescribed interval.
@@ 3 Theoretical guarantees
$f(\boldsymbol{y})=u_{v}(T,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ (19)
for a given node $v$, where $\boldsymbol{u}(\cdot,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ is the unique solution to the problem (4) with $\boldsymbol{C}(t)=\boldsymbol{C}(t,\boldsymbol{y})$. As illustrated in Section 2.4, functions that can be extended in a holomorphic way have desirable properties with respect to the recoverability of their polynomial approximation; hence our goal is to show that there exists a holomorphic extension for $f$. We will use [23] and [1, Lemma 4.3] which states the following:
@@
$\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert\leq L\lVert\boldsymbol{u}-\boldsymbol{v}\rVert,\quad\forall t,s\in[0,T],\quad\forall\boldsymbol{u},\boldsymbol{v}\in\mathbb{C}^{n},\forall\boldsymbol{z}\in\mathcal{K},$ (20)
for some norm $\lVert\cdot\rVert$ over $\mathbb{C}^{n}$. Then the Volterra integral equation
$\boldsymbol{u}(t,\boldsymbol{z})=\boldsymbol{g}(t,\boldsymbol{z})+\int_{0}^{t}\boldsymbol{h}(t,s,\boldsymbol{u}(s,\boldsymbol{z}),\boldsymbol{z})ds,\quad\forall t\in[0,T],$
admits a unique solution $t\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ for every $\boldsymbol{z}\in\mathcal{K}$ and the mapping $(t,\boldsymbol{z})\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ is continuous in $0,T]\times\mathcal{K}$. In addition, let $\mathring{\mathcal{K}}$ be the interior of $\mathcal{K}$ and assume that $\boldsymbol{z}\mapsto\boldsymbol{g}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for any fixed $t\in[0,T]$ and $(\boldsymbol{v},\boldsymbol{z})\mapsto\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})$ is holomorphic in $\mathbb{C}^{n}\times\mathring{\mathcal{K}}$ for any fixed $(t,s)\in[0,T]^{2}$. Then $\boldsymbol{z}\mapsto\boldsymbol{u}(t,\boldsymbol{z})$ is holomorphic in $\mathring{\mathcal{K}}$ for any fixed $t\in[0,T]$.
@@
Theorem 1 (Holomorphy of parametric graph diffusion).
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2 for some compact $\mathcal{K}\subseteq\mathbb{C}^{d}$ such that $\mathcal{U}\subset\mathcal{K}$. Then the map $\boldsymbol{f}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}^{|V|}$ defined by $\boldsymbol{f}(\boldsymbol{y})=\boldsymbol{u(}T,\boldsymbol{y})$ satisfying problem (4) admits a holomorphic extension $\tilde{\boldsymbol{f}}$ to $\mathring{\mathcal{K}}$. Moreover, for any compact subset $\mathcal{H}\subset\mathring{\mathcal{K}}$, we have
$\|\tilde{f}\|_{L^{\infty}(\mathcal{H)}}\leq B(\boldsymbol{u}_{0},T,\boldsymbol{M}),$ (21)
@@
$B(\boldsymbol{u}_{0},T,\boldsymbol{M})=\|\boldsymbol{u}_{0}\|_{2}\cdot\exp\left(2\cdot\int_{0}^{T}\sup_{\boldsymbol{z}\in\mathcal{H}}\|\boldsymbol{M}(t,\boldsymbol{z})\|_{2\to 2}\,dt\right)<\infty$ (22)
@@
We define a holomorphic extension $\tilde{\boldsymbol{f}}$ in a natural way using the complex-valued extension $\tilde{\boldsymbol{C}}$ to $\boldsymbol{C}$, whose existence is guaranteed by Assumption 2. Now, observe that we can reformulate problem (4) (where $\boldsymbol{C}$ is replaced by $\tilde{\boldsymbol{C}}$) as the parametric Volterra integral equation $\tilde{\boldsymbol{u}}(t,\boldsymbol{z})=\boldsymbol{u}_{0}+\int_{0}^{t}\boldsymbol{h}(t,s,\tilde{\boldsymbol{u}}(s,\boldsymbol{z}),\boldsymbol{z})ds$, $t\in[0,T],$ where $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$ and $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})=\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(t,\boldsymbol{z}),\boldsymbol{W}(t))$, with $\boldsymbol{D}$ defined as in equation (3).
@@
  • $\boldsymbol{g}(t,\boldsymbol{z})\equiv\boldsymbol{u}_{0}$ is continuous in $0,T]\times\mathcal{K}$, and $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$ is continuous in $0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$;
@@
Firstly, $\boldsymbol{g}(t,\boldsymbol{z})\equiv\boldsymbol{u}_{0}$ is a constant function, therefore $\boldsymbol{g}$ is continuous in $0,T]\times\mathcal{K}$ and holomorphic in $\mathring{\mathcal{K}}$ for all fixed $t\in[0,T]$.